# Formal Analysis of the Contract Automata Runtime Environment with Uppaal: Modelling, Verification and Testing

### v2

@@ 8. Conclusion and Future Work
At present, the different artifacts such as the model, source code, tests, and tracing information are manually kept aligned. This process demands substantial effort whenever a new version of CARE is introduced, as each artifact needs to be updated accordingly. Future work involves studying techniques for automatic alignment of these artifacts. Furthermore, it would be interesting to exploit the facilities provided by the recently introduced tool Uppex [59] to factorize the different variants of the model discussed in the paper under a single, configurable Uppex model, to automate the selection of a configuration and more generally to facilitate the collaboration between the software developer and the Uppaal modeller, whenever they are different persons.

### v3

@@ 8. Conclusion and Future Work
At present, the different artifacts such as the model, source code, tests, and tracing information are manually kept aligned. This process demands substantial effort whenever a new version of CARE is introduced, as each artifact needs to be updated accordingly. Future work involves studying techniques for automatic alignment of these artifacts. Furthermore, it would be interesting to exploit the facilities provided by the recently introduced tool Uppex [PPN+23] to factorize the different variants of the model discussed in the paper under a single, configurable Uppex model, to automate the selection of a configuration and more generally to facilitate the collaboration between the software developer and the Uppaal modeller, whenever they are different persons.